The following information is provided about an open economy with a government. Use the
information to answer the questions that follow:

C = 450 + 0.4Y
I = 350
G = 150
X = 70
Z = 35 + 0.1Y
T = 0.15Y
Yf = 1550

Q.2.1 Calculate the level of autonomous spending in this economy.
1
Expert's answer
2020-11-30T16:29:43-0500

Consumption: C = 450 + 0.4Y where Y is income

Investment: I = 350

Government spending: G = 150

Exports: X = 70

Imports: Z = 35 + 0.1Y

Taxes: T = 0.15Y

Full employment income: Yf = 1550

Y = C + I + G + X – Z

Y = 450 + 0.4Y + 350 + 150 + 70 – 35 – 0.1Y

Y = 985 + 0.4(Y – T) – 0.1Y

Y = 985 + 0.4(Y – 0.15Y) – 0.1Y

Y = 985 + 0.24Y

Autonomous Spending = 985